OpenStudy (anonymous):

6√72x^2 ,would someone please explain to me in detail how i would solve this

OpenStudy (anonymous):

solve? please provide an equation

OpenStudy (anonymous):

o i'm sorry, i meant factor

OpenStudy (anonymous):

\[6\sqrt{72x ^{2}}\] is this what you have?

OpenStudy (anonymous):

Yes sir, you are correct

OpenStudy (mertsj):

\[6\sqrt{72x^2}=6\sqrt{36x^2(2)}=6\sqrt{36}\sqrt{x^2}\sqrt{2}\]

OpenStudy (mertsj):

Now, you can finish it

OpenStudy (anonymous):

I see, but can't you continue to factor out the 36 :o

OpenStudy (anonymous):

\[6\sqrt{36\times2\times x^{2}}=6\times \sqrt{36}\times \sqrt{x ^{2}}\times \sqrt{2}=36x \sqrt{2}\]

OpenStudy (anonymous):

Thank you very much, i did the math and i recieved the correct answer ^.^